### v3

# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

Abstract
This paper investigates the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium (NE) in potential games. In such games, an efficient NE is defined as the maximizer of the potential function. Previous literature provides asymptotic convergence rates to efficient Nash equilibria, and existing finite-time rates are limited to potential games with further assumptions such as the interchangeability of players. In this paper, we prove the first finite-time convergence to an $\epsilon$-efficient NE in general potential games. Our bounds depend polynomially on $1/\epsilon$, an improvement over previous bounds that are exponential in $1/\epsilon$ and only hold for subclasses of potential games. We then strengthen our convergence result in two directions: first, we show that a variant of log-linear learning that requires a factor $A$ less feedback on the utility per round enjoys a similar convergence time; second, we demonstrate the robustness of our convergence guarantee if log-linear learning is subject to small perturbations such as alterations in the learning rule or noise-corrupted utilities.
keywords:
Game theory, potential games, log-linear learning, efficient Nash equilibrium.
1 Introduction
Interactions of multiple agents are at the heart of many applications in transportation networks, auctions, telecommunication networks, and multi-robot systems. In game theory, the Nash equilibrium is a popular solution concept to describe outcomes of a multi-agent system (Nash, 1951). Fundamental considerations in game theory are whether an NE exists, if strategic players can learn it, and if so at which speed they can learn it. Furthermore, for practical purposes it is important to understand which NE is learned. This is particularly pertinent in games that admit a social welfare function, as it enables the definition of an efficient NE as an NE that maximizes social welfare. The social welfare is typically an aggregate measure of individual utilities such as their sum or a measure of fairness. In distributed control of engineering systems, for example, studying efficient NEs is important, as the aim is to optimize some global objective function in a distributed manner.
The class of potential games (Monderer and Shapley, 1996b) lends itself to studying efficient NEs since every joint action maximizing the potential corresponds to an NE. This property follows from the fact that in potential games, the difference in a player’s utility generated by a unilateral change of her action equals the difference in potential. Consequently, if the social welfare is aligned with the potential function, meaning that an increase in social welfare is associated with an increase in potential (Paccagnan et al., 2022), then any joint action maximizing the potential is an efficient NE (Marden and Shamma, 2015). In identical interest games, for example, maximizing the aggregated utility is achieved by maximizing the potential function which is trivially given by the common utility function. In the example of coverage problems, the goal is to assign players to locations to achieve maximal coverage. When each player’s utility is designed as the marginal contribution then this goal is achieved by maximizing the resulting potential function (Marden and Wierman, 2013).
In this paper, we study the speed of convergence to approximate efficient NE in potential games.
1.1 Related work
In potential games, many learning rules are proven to converge to an NE, such as iterative best-response dynamics (Rosenthal, 1973; Awerbuch et al., 2008; Chien and Sinclair, 2011), no-regret algorithms (Krichene et al., 2015; Palaiopanos et al., 2017; Heliou et al., 2017), and fictitious play (Monderer and Shapley, 1996a, b). However, only log-linear learning (Blume, 1993; Young, 1993) is known to converge to an efficient Nash equilibrium. [^1] In log-linear learning, players asynchronously choose actions with probability proportional to their exponentiated utility. Due to this desirable property, variants of log-linear learning were proposed in subsequent works (Marden et al., 2007; Marden and Shamma, 2012). For example, Leslie and Marden (2011) proves that log-linear learning also handles the practical setting where the observed utilities are corrupted by noise. Moreover, Arslan et al. (2007) proposes binary log-linear learning, a slight modification of log-linear learning that only requires two points of feedback per round. Furthermore, (Tatarenko, 2017) provides an asymptotic rate of convergence to an efficient NE when the parameter controlling the players’ rationality in log-linear learning is time-varying. Another line of research studies how fast log-linear learning converges to its stationary distribution (Auletta et al., 2010, 2011, 2012) and how it behaves before reaching the stationary distribution (Auletta et al., 2012), but do not focus on Nash equilibria.
Few past works provide finite-time guarantees for log-linear learning to an $\epsilon$-efficient NE, an action profile whose potential is $\epsilon$-close to the maximum value. For instance, for atomic routing games with polynomial costs of degree at most $p$, Asadpour and Saberi (2009) proves a convergence time exponential in $p$ and $1/\epsilon$ and polynomial in $N$, where $N$ denotes the number of players. Moreover, Montanari and Saberi (2008, 2010) study games with graph structures between players and prove an exponential convergence time in $N$ and $1/\epsilon$ in the worst case. Finally, in potential games with interchangeable players and a Lipschitz-continuous potential function, Shah and Shin (2010) shows a convergence time exponential in $A$ and $1/\epsilon$ and linear in $N$, where $A$ is the number of actions per player. The latter result was extended in (Borowski and Marden, 2015) to semi-anonymous potential games, which are potential games with groups of interchangeable players.
Although finite-time convergence of log-linear learning to an $\epsilon$-efficient NE was established for different sub-classes of potential games, such guarantees do not exist for general potential games. In this paper, we focus on deriving the first finite-time convergence guarantees of log-linear learning that hold for general potential games; are polynomial in $1/\epsilon$; and remain valid under relaxed assumptions such as limited access to feedback and noisy utilities. [^2] For general potential games, an exponential dependence on $N$ in the convergence time is unavoidable since finding an $\epsilon$-efficient NE is equivalent to computing an $\epsilon$-optimizer of the potential function, which is NP-complete (Burer and Letchford, 2012).
1.2 Contributions
We study the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium in general potential games. Our contributions are summarized as follows:
  • We prove a convergence time of $\tilde{\mathcal{O}}((A^{N}/\epsilon)^{\frac{1}{\max\{\epsilon,\Delta\}}})$ to an $\epsilon$-efficient NE (Theorem 3.1), where $\Delta$ is a problem-dependent constant. If in addition, the players are interchangeable, then an $\epsilon$-NE is reached in $\tilde{\mathcal{O}}((\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$ which in contrast to general potential games is polynomial in $N$ as well (Corollary B.7).
  • We consider two variants, binary log-linear learning (Theorem 4.1) and perturbed log-linear learning (Theorem 4.3) motivated by limited feedback and noise corrupted utilities, respectively. For these variants, we show convergence guarantees that are polynomial in $1/\epsilon$.
On the technical side, past works (Blume, 1993; Young, 1993; Marden and Shamma, 2012; Shah and Shin, 2010) established asymptotic convergence guarantees by leveraging that log-linear learning induces a Markov chain. To obtain our novel finite-time results we build on this connection and develop new Markov chain results that can be summarized as follows:
  • We derive a mixing time bound for a class of Markov chains which includes those induced by log-linear learning and binary-log linear learning. It follows from an adaptation of arguments in (Diaconis and Saloff-Coste, 1996) (Lemma 2.2) and a novel bound on the log-Sobolev constant (Lemma 3.4) that we derive for this class of Markov chains.
  • We derive a tight Lipschitz constant for the known result regarding the Lipschitz-continuity of stationary distributions of Markov chains as a function of their transition matrix (Lemma 4.2). We leverage this result to study the convergence of learning rules such as perturbed log-linear learning for which the explicit stationary distributions are unknown (Theorem 4.3).
Notations
We denote by $[N]$ the set $\{1,\ldots,N\}$. For a finite set $\mathcal{X}$, we denote by $\Delta(\mathcal{X})$ the probability simplex over $\mathcal{X}$, and by $\mathds{1}_{a\in\mathcal{X}}$ the indicator function of $\mathcal{X}$. Finally, we use the big-$\mathcal{O}$ notations $\tilde{\mathcal{O}}$ and $\tilde{\Omega}$ to hide logarithmic terms.
2 Preliminaries
2.1 Problem setup
We consider a repeated potential game with $N$ players. Every player has an action set $\mathcal{A}$ of cardinality $A<\infty$, which for simplicity we assume to be the same for all players. The utility of player $i$ is a mapping $U_{i}:\mathcal{A}^{N}\rightarrow[0,1]$ from the joint action space $\mathcal{A}^{N}$ to $[0,1]$. In a potential game, the utility functions are characterized by a potential function $\Phi:\mathcal{A}^{N}\rightarrow\mathbb{R}$ such that the following holds:
$U_{i}(a_{i},a_{-i})-U_{i}(a_{i}^{\prime},a_{-i})=\Phi(a_{i},a_{-i})-\Phi(a_{i}^{\prime},a_{-i}),\quad\forall i\in[N],\forall a_{i},a_{i}^{\prime}\in\mathcal{A},\forall a_{-i}\in\mathcal{A}^{N-1}.$
In this paper, we assume that an initial action profile is drawn from a distribution $\mu^{0}$. The potential game is then repeated over multiple rounds. We assume that at each round $t$, a player $i\in[N]$ is uniformly chosen and selects her action based on some learning rule while the other players remain with their previous action, i.e., $a_{-i}^{t}=a_{-i}^{t-1}$.
A common assumption in game theory is that the players are rational, i.e., that they seek to maximize their utility. Under this assumption, a natural solution concept is the pure Nash equilibrium (Nash Jr, 1950).
Definition 2.1 (Nash equilibrium).
A pure Nash equilibrium is an action profile $(a_{i}^{NE})_{i\in[N]}\in\mathcal{A}^{N}$ such that every player is playing a best response to the other players’ actions $a_{-i}^{NE}$, i.e.,
$U_{i}(a_{i},a_{-i}^{NE})\leq U_{i}(a_{i}^{NE},a_{-i}^{NE}),\quad\forall i\in[N],\forall a_{i}\in\mathcal{A},$
where we define $a_{-i}:=(a_{j})_{j\in[N]\setminus\{i\}}$ as the joint action of all players except $i$.
In a pure NE, no player can improve her utility by unilaterally changing her action. Hereafter, we simply refer to a pure Nash equilibrium as a Nash equilibrium.
In potential games, several Nash equilibria may exist as shown in the two-player potential game example below. Here, action profiles $(A,A)$ and $(B,B)$ are both Nash equilibria. However, the value of the potential function may differ for different Nash equilibria, as is the case in the example.
$\bordermatrix{&A&B&\cr A&(5,2)&(-1,-2)\!\!\!\!\!\cr B&(-5,-4)&(1,4)\!\!\!\!\!\!},\quad\text{with the potential:}\quad\bordermatrix{&A&B&\cr A&4&0\!\!\!\!\!\!\!\!\cr B&-6&2\!\!\!\!\!\!\!\!}.$
[…]
In this work, we are interested in learning rules that converge, in expectation, to an $\epsilon$-efficient NE, i.e., an action profile maximizing the potential up to an additive constant:
$\mathbb{E}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon,$
where the randomness stems from the distribution of action profiles at round $t$. The number of rounds needed to find an $\epsilon$-efficient NE denotes the convergence time.
Connection to Markov chains:
If all players apply a learning rule that relies exclusively on the utility of the current action profile of the other players then the learning dynamics induce a Markov chain over the state space $\mathcal{A}^{N}$. In particular, the state of the Markov chain at time $t$ corresponds to the action profile at round $t$ and the learning rule of each player specifies the transition dynamics. It follows that we can analyze the convergence time of log-linear learning by studying the associated Markov chain. We, therefore, provide relevant concepts for Markov chains in the next section.
2.2 Background on Markov chains
We briefly review concepts and properties of Markov chains used throughout this paper. Consider a time-homogeneous Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ over the state space $\mathcal{A}^{N}$ with a transition matrix $P\in\mathbb{R}^{{A^{N}}\times A^{N}}$. The ergodic theorem (Levin and Peres, 2017) states that if a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ is irreducible and aperiodic, then it has a unique stationary distribution $\mu$, and from any initial distribution $\mu^{0}$ the distribution $\mu^{t}=\mu^{0}P^{t}$ converges to $\mu$. The convergence time to the stationary distribution is quantified by the mixing time:
$t_{\operatorname{mix}}^{P}(\epsilon):=\min\{t\in\mathbb{N}\>|\>\|\mu^{t}-\mu\|_{TV}\leq\epsilon\},$ (1)
where the total variation distance is defined as $\|\mu^{t}-\mu\|_{TV}:=\frac{1}{2}\sum_{a\in\mathcal{A}^{N}}|\mu^{t}(a)-\mu(a)|$. Based on a remark in (Diaconis and Saloff-Coste, 1996, Section 3) we now derive a bound on the mixing time of Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$.
Lemma 2.2.
If $P$ is irreducible and aperiodic, then the mixing time has the following upper bound:
$t_{\operatorname{mix}}^{P}(\epsilon)\leq\frac{1}{\rho(PP^{*})}\left(\log\log\frac{1}{\mu_{\min}}+2\log\frac{1}{\epsilon}\right),$ (2)
[…]
We briefly discuss the mixing time bound above and provide a proof and a thorough discussion in Appendix A. While classical approaches commonly bound the mixing time by the spectral gap defined in Equation (11) in Appendix A, bounds using log-Sobolev constants are often significantly tighter. Indeed, the mixing time upper bound using the log-Sobolev constant grows as $\mathcal{O}(\log\log(1/\mu_{\min}))$ whereas bounds using the spectral gap grow as $\mathcal{O}(\log(1/\sqrt{\mu_{\min}}))$. However, unlike the spectral gap deriving log-Sobolev constants can be extremely difficult. So constants have not been well-explored. In the next section in Lemma 3.4, we derive a novel bound on the log-Sobolev constant of a class of Markov chains.
3 Convergence of log-linear learning
In this section, we review the well-established log-linear learning rule (Blume, 1993) and state our main theoretical result on the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium. Our convergence analysis relies on the mixing time bound from Lemma 2.2 of the previous section.
3.1 Algorithm and background
We assume that all players follow the log-linear learning rule which is repeated over several rounds. At round $t$ a player denoted by $i$ is randomly chosen among all players and allowed to alter her action while the other players repeat their current action, i.e., $a_{-i}^{t}=a_{-i}^{t-1}$. Given full-information feedback, player $i$ observes her utility for all actions $a_{i}\in\mathcal{A}$ given the other players’ actions $a_{-i}^{t-1}$. Then, player $i$ samples an action from her strategy $p_{i}^{t}\in\Delta(\mathcal{A})$ such that:
$p_{i}^{t}(a_{i})=\frac{e^{\beta U_{i}(a_{i},a_{-i}^{t-1})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},a_{-i}^{t-1})}},\quad\forall a_{i}\in\mathcal{A},$ (3)
[…]
Log-linear learning induces an irreducible and aperiodic Markov chain $\{X_{t}\}_{t\in\mathbb{Z}_{+}}$ with a time-reversible transition matrix $P\in\mathbb{R}^{\mathcal{A}\times\mathcal{A}}$ (Marden and Shamma, 2012) given by:
$P_{a,\tilde{a}}=\frac{1}{N}\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}_{i}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)},$ (4)
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ of log-linear learning is given by (Blume, 1993):
$\mu(a)=\frac{e^{\beta\Phi(a)}}{\sum_{\tilde{a}\in A^{N}}e^{\beta\Phi(\tilde{a})}}\quad\forall a\in\mathcal{A}^{N}.$ (5)
The above can be verified by checking the detailed balance equations given by $\mu(a)P_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$. For $\beta\rightarrow\infty$, sampling an action profile $a\in\mathcal{A}^{N}$ from the stationary distribution $\mu$ returns a maximizer of the potential $\Phi(\cdot)$ with arbitrarily high probability. Thus, when all players adhere to log-linear learning with sufficiently large $\beta$, the global outcome, in the long run, will correspond to a potential maximizer, i.e., an efficient Nash equilibrium.
For sufficiently large $\beta$, it was shown that log-linear learning converges asymptotically to a potential maximizer and thus to an efficient Nash equilibrium (Blume, 1993; Young, 1993; Marden and Shamma, 2012). Except for a few works (Montanari and Saberi, 2010; Asadpour and Saberi, 2009; Shah and Shin, 2010), none of the previous works, however, have finite-time convergence guarantees to such an efficient Nash equilibrium, and (Montanari and Saberi, 2010; Asadpour and Saberi, 2009; Shah and Shin, 2010) have additional assumptions on the potential game. Thus, in the following section, we establish a bound on the convergence time of log-linear learning in general potential games.
3.2 Convergence time
We now state our main result on the convergence time of log-linear learning to an $\epsilon$-efficient NE. Before stating the result we briefly introduce some notation. Denote by $a^{*}$ a potential maximizer, i.e., $a^{*}\in\arg\max_{a\in\mathcal{A}^{N}}\Phi(a)$. The set of $\epsilon$-optimal action profiles is defined as $\mathcal{A}^{N}(\epsilon):=\{a\in\mathcal{A}^{N}|\Phi(a)\geq\Phi(a^{*})-\epsilon\}$ with cardinality $A^{N}(\epsilon)=|\mathcal{A}^{N}(\epsilon)|$. Furthermore, the suboptimality gap is defined as $\Delta:=\min_{a\in\mathcal{A}^{N}:\Phi(a)<\Phi(a^{*})}\left(\Phi(a^{*})-\Phi(a)\right)$ and is non-negative.
Theorem 3.1.
[…]
$\beta$ $\geq\frac{1}{\max\{\epsilon/2,\Delta\}}\log\left((A^{N}-A^{N}(\epsilon/2))\left(\frac{4}{\epsilon A^{N}(\epsilon/2)}-\frac{1}{A^{N}(\epsilon/2)}\right)\right).$ (6)
Then,
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon,$
for
$t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}.$
In other words, after running $t=\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{1/\max\{\epsilon,\Delta\}})$ rounds of log-linear learning with $\beta=\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon}$ the expected value of the potential of the joint action at time $t$ is $\epsilon$-optimal.
Theorem 3.1 provides the first finite-time convergence rates to an $\epsilon$-efficient NE in general potential games. In order to be $\epsilon$-close to an efficient NE $\beta$ must scale as $\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon})$. Furthermore, since the convergence time scales as $e^{4\beta}$ it grows polynomially in $\mathcal{A}$ and $1/\epsilon$ and exponentially in $N$. The exponential dependence on $N$ is unavoidable without further assumptions on the game. However, by deriving problem-dependent bounds we are the first to avoid the exponential dependence in $1/\epsilon$, see Table 1. Note that the case $\Delta=0$ is trivial as all action profiles are efficient NEs. In Appendix E, we conduct numerical experiments to showcase the performance of log-linear learning with $\beta$ specified as in Equation (6). To summarize, our experiments demonstrate that log-linear learning with such a $\beta$ converges to an $\epsilon$-efficient NE. In particular, it can be seen that the suboptimality gap $\Delta$ is an important feature to quantify the convergence time and that the convergence time increases polynomially in $1/\epsilon$ for $\epsilon<\Delta$.
{collect*}
mytable
Table 1: Convergence time of log-linear learning to an $\epsilon$-efficient NE.
Game setting Assumptions Convergence time
Routing game with $K$ vertices Cost functions of $\tilde{\mathcal{O}}(e^{\frac{N}{\epsilon}})$
(Asadpour and Saberi, 2009) degree at most $p$
Potential game with interchangeable $\lambda$-Lipschitz continuous $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$
players (Shah and Shin, 2010) potential function
Potential game with interchangeable $A\geq 4$ $\tilde{\mathcal{O}}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
players (Corollary B.7)
Potential games (Theorem 3.1) $A\geq 4$ $\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
Proof 3.2.
outline Here, we provide an outline of the proof, the full proof is deferred to Appendix B.1. The argument begins with the following decomposition:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\underbrace{\mathbb{E}_{a\sim\mu}[\Phi(a)]}_{\text{First term}}-2\underbrace{\|\mu^{t}-\mu\|_{TV}}_{\text{Second term}}\underbrace{\max_{a\in\mathcal{A}^{N}}\Phi(a)}_{\leq 1}$
First term: To control the first term above, we rely on the novel lemma below which shows that an action profile sampled from $\mu$ is in expectation $\epsilon/2$-optimal if $\beta$ is sufficiently large.
Lemma 3.3.
For any $\epsilon\in(0,1)$, if all players adhere to log-linear learning with:
$\beta$ $\geq\frac{1}{\max\{\epsilon,\Delta\}}\log\left((A^{N}-A^{N}(\epsilon))\left(\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}\right)\right),$
then it holds that $\mathbb{E}_{a\sim\mu}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon.$
The proof of this lemma is provided in Appendix B.2. We use it to control the first term as:
$\mathbb{E}_{a\sim\mu}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon/2$
when $\beta$ is defined as in Equation (6).
Second term: We can control the mixing time of log-linear learning using the bound of Lemma 2.2:
$\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4$
for $t\geq\frac{1}{\rho(PP^{*})}(\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon})$. We then bound $\rho(PP^{*})$ using the novel lemma below.
Lemma 3.4.
Consider a Markov chain over state space $\mathcal{A}^{N}$ with $A\geq 4$. Assume that there exists $p_{\min},p_{\max}\in(0,1]$, such that the corresponding transition matrix $P\in\mathbb{R}^{A^{N}\times A^{N}}$ satisfies: [^5] Such a $p_{\min}\in(0,1]$ exists if the probability of transitioning from any state to a neighboring state is positive.
$\frac{1}{N}p_{\min}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}\leq P_{a,\tilde{a}}\leq\min\{1,\frac{1}{N}p_{\max}\}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$ (7)
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. Then the log-Sobolev constant $\rho(PP^{*})$ of matrix $PP^{*}$ is lower bounded by:
$\rho(PP^{*})\geq\frac{16\pi^{2}A^{N-2}\mu_{\min}p_{\min}^{3}}{25N^{2}},$
where $\mu$ is the stationary distribution of the Markov chain induced by $P$ and $\mu_{\min}=\min_{a\in\mathcal{A}^{N}}\mu(a)$.
The proof of this lemma is provided in Appendix B.3. It provides a bound on $\rho(PP^{*})$ for any transition matrix that satisfies Equation (7). It is in particular applicable to the Markov chain induced by log-linear learning since the transition matrix specified in Equation (4) satisfies Equation (7).
Combination: Combining the two parts of the proof we deduce that:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\mathbb{E}_{a\sim\mu}[\Phi(a)]-2\|\mu^{t}-\mu\|_{TV}\max_{a\in\mathcal{A}^{N}}\Phi(a)$
$\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\frac{\epsilon}{2}-\frac{2\epsilon}{4},$
for $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)$. This concludes the proof of the theorem.
In the following, we additionally assume that the potential game is symmetric, i.e., players are interchangeable. [^6] A definition of a symmetric game is given in Appendix B.4, see Definition B.6. In these games, a player’s utility depends solely on the number of players selecting each action and not on their identity. Technically speaking, in this case, log-linear learning induces a Markov chain on a modified state space of cardinality $\mathcal{O}(N^{A})$, see Equation (22), rather than on the state space $A^{N}$. Exploiting the smaller state space, we derive a bound on the convergence time of log-linear learning that depends polynomially not only on $1/\epsilon$ but also on $N$.
Corollary 3.5 (Sketch).
Consider a symmetric potential game. For any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to modified log-linear learning (Shah and Shin, 2010) with $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log(\frac{N^{A}}{\epsilon}))$. Then, $\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon$ for $t=\tilde{\mathcal{O}}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$.
We provide a full statement of the corollary and its proof in Appendix B.4. For symmetric potential games with a $\lambda$-Lipschitz-continuous potential function, Shah and Shin (2010) proves a convergence time of $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$. In comparison, relaxing the Lipschitzness assumption comes at a small cost of polynomial dependence on $N$ rather than a linear one. However, our result greatly improves the dependence on $\epsilon$ to a polynomial one.
4 Robustness of log-linear learning
In this section, we show that our convergence time guarantee is robust to settings where players have less feedback information and to settings where log-linear learning is subject to perturbations such as noisy utility observations.
4.1 Reduced feedback
Log-linear learning requires players to access their utilities for all possible actions given the other players’ actions. Having such full-information feedback when action sets are large can be demanding. Binary log-linear learning (Arslan et al., 2007; Marden et al., 2007) alleviates this limitation by requiring two-point feedback, reducing the feedback needed by a factor $A$ per round. Now, we briefly review the binary-log-linear learning rule. Then, we derive the first finite-time convergence bound of binary log-linear learning to an $\epsilon$-efficient Nash equilibrium, showing that the deterioration in convergence time is a constant.
Binary log-linear learning proceeds as log-linear learning with the distinction that the player $i$ allowed to alter her action first samples a trial action $\tilde{a}_{i}$ uniformly from her action set $\mathcal{A}$. She then plays according to the strategy:
$p_{i}^{t}(a)=\begin{cases}\frac{e^{\beta U_{i}(a,a_{-i}^{t-1})}}{e^{\beta U_{i}(a_{i}^{t-1},a_{-i}^{t-1})}+e^{\beta U_{i}(\tilde{a}_{i},a_{-i}^{t-1})}}&\text{ for }a\in\{a_{i}^{t-1},\tilde{a}_{i}\}\\ 0&\text{ otherwise}\end{cases}$
Here, player $i$ can either repeat her action $a_{i}^{t-1}$ or play one other randomly sampled action $\tilde{a}_{i}$ rather than any action $a_{i}\in\mathcal{A}$ as in log-linear learning.
Theorem 4.1.
Consider a potential game with potential function $\Phi:\mathcal{A}^{N}\rightarrow[0,1]$ and $A\geq 4$. For any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to binary log-linear learning with $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{\mathcal{A}^{N}}{\epsilon}).$ Then, it holds that $\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon$ for
$t\geq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)\approx\tilde{\mathcal{O}}\left(N^{2}A^{5}\left(\frac{A^{N}}{\epsilon}\right)^{\frac{N}{\max\{\epsilon,\Delta\}}}\right).$
Theorem 4.1 shows the first finite-time convergence guarantee for binary log-linear learning, we provide its detailed proof in Appendix C. It is remarkable that with significantly less feedback per round, binary log-linear achieves the same convergence speed as log-linear learning up to a factor of $8$ (Theorem 3.1). This raises the question of whether two-point feedback is sufficient for learning $\epsilon$-efficient Nash equilibria or whether the convergence time bounds we proved for log-linear learning are loose.
4.2 Perturbed log-linear learning
Classical log-linear learning relies on two limiting assumptions: 1) Players have access to their exact utilities. However, in real-world applications, the presence of noise is typical as uncertainties and hidden factors generate inexact measurements. 2) Players are rational. However, empirical evidence suggests that players have limited rationality and therefore may occasionally deviate from the log-linear learning rule in practical scenarios. Our next result generalizes Theorem 3.1 to the case where the log-linear learning rule is subject to small perturbations. As we will show, this generalization can address the two limitations above.
We first derive a general statement for Markov chains which shows that the induced stationary distribution is Lipschitz-continuous as a function of the transition matrix. We then use this lemma to prove our main result on the convergence time of perturbed log-linear learning to an $\epsilon$-efficient NE.
Lemma 4.2 (Lipschitzness).
Consider two irreducible and aperiodic transition matrices $P_{1},P_{2}\in\mathbb{R}^{A^{N}\times A^{N}}$. Let $\mu_{1}$ and $\mu_{2}$ be the stationary distributions of the Markov chains induced by $P_{1}$ and $P_{2}$, respectively. Then, the following holds:
$\|\mu_{1}-\mu_{2}\|_{2}\leq\min\{L(P_{1}),L(P_{2})\}\|P_{1}-P_{2}\|_{2},$
where $L(P_{k}):=\frac{2A^{N}}{\rho(P_{k}P_{k}^{*})}(\log\log\frac{1}{\mu_{k,\min}}+\log(8A^{N}))$ and $\mu_{k,\min}=\min_{a\in\mathcal{A}^{N}}\mu_{k}(a)$ for $k=1,2$.
We provide a proof in Appendix D.4. Compared to the result of (Zhang et al., 2023), we considerably improve the Lipschitz constant by using the mixing time bound based on log-Sobolev inequalities (Lemmas 2.2 and 3.4). In particular, (Zhang et al., 2023, Lemma 24) entails a Lipschitz constant $L=\tilde{\mathcal{O}}((e/p_{\min})^{N})$ while Lemma 4.2 implies that $L=\tilde{\mathcal{O}}(1/(\mu_{\min}p_{\min}^{3}))$. [^7] This can be seen by injecting the log-Sobolev bound of lemma 3.4 into our Lemma above. Now, we state the main result of this section.
Theorem 4.3.
Consider a potential game with a potential function $\Phi:\mathcal{A}^{N}\rightarrow[0,1]$ and $A\geq 4$. Furthermore, consider a learning rule with transition matrix $P$ and assume that there exists $p_{\min},p_{\max}\in(0,1]$, such that for all $a,\tilde{a}\in\mathcal{A}$ it holds that:
$\frac{1}{N}p_{\min}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}\leq P_{a,\tilde{a}}\leq\min\{1,\frac{1}{N}p_{\max}\}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}.$ (8)
For any $\epsilon\in(0,1)$ and any initial distribution $\mu^{0}$, assume all players adhere to this learning rule with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$. Then,
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}}\Phi(a)-\epsilon-L\sqrt{A^{N}}\|P-P_{\ell}\|_{2},$
for
$t\geq\frac{25N^{3/2}e^{N}}{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}\log\left(\frac{4A^{N}}{\epsilon^{2}}\log\frac{e^{N}}{p_{\min}^{N}\sqrt{2\pi N}}\right),$
where $P_{\ell}$ is the transition matrix of log-linear learning and $L$ is a Lipschitz constant of order $\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
Theorem 4.3 shows that small perturbations of the log-linear learning rule do not compromise the convergence to an $\epsilon$-efficient NE. In particular, if the players follow a learning rule $P$ with $\|P-P_{\ell}\|_{2}=\mathcal{O}(\epsilon/(L\sqrt{A^{N}}))$ then they converge to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$. On the other hand, due to the unavailability of the stationary distribution of the perturbed learning rule, we suffer an extra factor of $(N/p_{\min})^{N}/N!$ in the convergence time guarantee compared to log-linear learning.
We now consider two explicit types of perturbations: Noisy utilities and a modified learning rule.
4.2.1 Corrupted utilities with additive noise
In the following, we assume that players observe noise-corrupted utilities $(\hat{U}_{i})_{i\in[N]}$ which satisfy:
$\hat{U}_{i}(a_{i},a_{-i})=U_{i}(a_{i},a_{-i})+\xi_{i}(a_{i},a_{-i}),\quad\forall(a_{i},a_{-i})\in\mathcal{A}^{N}$ (9)
where $\xi_{i}(a_{i},a_{-i})\in[-\xi,\xi]$ is a bounded noise term. Alternatively, the noise could be assumed to be centered i.i.d. random variables with bounded variance (Leslie and Marden, 2011). Using Theorem 4.3, we hereafter prove that log-linear learning is robust to noisy feedback.
Corollary 4.4.
Consider the setting of Theorem 4.3 with noise-corrupted utilities as in Equation (9). If all players adhere to log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$ and $\xi\leq 1/(2\beta)$, then
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}}\Phi(a)-\epsilon-\frac{7LA^{3N/2}}{2N}\beta\xi,$
for $t=\mathcal{O}\left(N^{3/2}A^{3}e^{N+\beta(1+2\xi)(N+3)}\log\frac{1}{\epsilon^{2}}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
The proof is provided in Appendix D.2. Corollary 4.4 shows that log-linear learning with corrupted utilities converges to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$ if the corruption magnitude $\xi$ is sufficiently small. Our finite-time convergence result extends previous works on robust learning which provide asymptotic guarantees (Leslie and Marden, 2011; Lim and Shamma, 2013; Bravo and Mertikopoulos, 2017). The key to this result lies in showing that the transition matrix of the Markov chain induced by corrupted utilities is close to its corruption-free counterpart.
4.2.2 Log-linear learning mixed with uniform exploration
In the following, we assume players occasionally explore actions randomly. A modification of log-linear learning based on the fixed-share algorithm (Herbster and Warmuth, 1998) can reflect such a random behavior. In the so-called fixed-share log-linear learning, a player $i$ is randomly chosen and allowed to alter her action. Player $i$ samples her new action from the following distribution:
$\hat{p}_{i}^{t}(a_{i})=\frac{\xi}{A}+\frac{(1-\xi)e^{\beta U_{i}(a_{i},a_{-i}^{t-1})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},a_{-i}^{t-1})}},\quad\forall a_{i}\in\mathcal{A}.$
The exploration parameter $\xi\in(0,1)$ determines how likely a player is to act randomly, where a value of $\xi=1$ corresponds to a uniform action sampling while $\xi=0$ corresponds to standard log-linear learning. For simplicity, we focus on the full-information case but fixed-share log-linear learning can easily be adapted to the binary setting. Note that this modification resembles the $\epsilon$-Hedge strategy (Heliou et al., 2017) in the expert advice literature, and under binary feedback, this modification resembles the Epx3.P strategy (Auer et al., 2002; Bubeck et al., 2012) in the multi-armed bandit literature. Here, the fixed share $\xi/A$ ensures a lower bound on the exploration.
Without explicit knowledge of the stationary distribution of this learning rule, we can apply Theorem 4.3 to deduce the following result.
Corollary 4.5.
Consider the setting of Theorem 4.3, where all players adhere to fixed-share log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$. Then, for any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$ we have:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}}\Phi(a)-\epsilon-\frac{LA^{N}}{\sqrt{N}}\xi,$
for $t=\mathcal{O}\left(N^{3/2}A^{N+3}e^{\beta(N+3)}/(1-\xi)^{N+3}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
We provide a proof in Appendix D.3. Corollary 4.5 guarantees the convergence of fixed-share log-linear learning to an $\epsilon$-efficient Nash equilibrium in time polynomial in $1/\epsilon$ if the exploration parameter $\xi$ is sufficiently small. The key to this result is to show that the transition matrix of fixed-share log-linear learning is close to the transition matrix of the unperturbed learning rule in terms of the $\ell_{2}$ distance.
5 Conclusion
We provided the first finite-time convergence guarantees to an $\epsilon$-efficient NE for potential games using a novel mixing-time bound based on a log-Sobolev constant. In particular, we guarantee a polynomial dependence on $1/\epsilon$ using a problem-dependent analysis. Furthermore, under the additional assumption that the game is symmetric, we showed that the exponential dependence on the number of players $N$ present in our bound on the convergence time can be avoided. To deal with reduced feedback, i.e., two-point feedback on the utility, we considered binary log-linear learning and showed that it enjoys the same convergence time as log-linear learning up to numerical constants. Finally, we proved that the convergence time of log-linear is not hindered by corruptions of the observed utilities by bounded noise or by small perturbations in the learning rule. The relevance of this result is twofold: First, our analysis does not rely on characterizing the stationary distribution of this perturbed Markov chain; Second, the presence of noise is ubiquitous in real-world applications and it is therefore crucial that the implemented learning rule is robust to such corruptions.
In future work, we are interested in providing lower bounds on the convergence time of log-linear learning to an $\epsilon$-efficient NE. Such bounds would allow us to assess the tightness of our results.
Appendix A Markov chains and mixing times
In the following, we define the log-Sobolev constant and the spectral gap of a Markov chain. We then discuss the implications of Lemma 2.2 and provide a proof of the lemma.
Definition A.1 ((Diaconis and Saloff-Coste, 1996)).
Consider a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ over state space $\mathcal{A}^{N}$ with transition matrix $P$ and stationary distribution $\mu$. The log-Sobolev constant $\rho(P)$ is defined as:
$\rho(P):=\inf_{\mathcal{L}_{\pi}(f^{2})\neq 0}\frac{\mathcal{E}_{P}(f,f)}{\mathcal{L}_{\pi}(f^{2})}.$ (10)
The spectral gap $\lambda(P)$ is defined as:
$\lambda(P):=\inf_{\text{Var}_{\pi}(f)\neq 0}\frac{\mathcal{E}_{P}(f,f)}{\text{Var}_{\pi}(f)}.$ (11)
For any $f:\mathcal{A}^{N}\rightarrow\mathbb{R}$, the Dirichlet form $\mathcal{E}_{P}(f,f)$ is defined by:
$\mathcal{E}_{P}(f,f)=\langle f,(I-P)f\rangle_{\pi}=\frac{1}{2}\sum_{a,\tilde{a}\in A^{N}}(f(a)-f(\tilde{a}))^{2}P_{a,\tilde{a}}\mu(a),$
the entropy-like quantity $\mathcal{L}(f^{2})$ is defined by:
$\mathcal{L}(f^{2})=\sum_{a\in A^{N}}f(a)^{2}\log\frac{f(a)^{2}}{\|f\|_{2}^{2}}\mu(a),$
and the variance $\text{Var}_{\pi}(f)$ is defined by:
$\text{Var}_{\pi}(f)=\sum_{a,\tilde{a}\in A^{N}}(f(a)-f(\tilde{a}))^{2}\mu(a)\mu(\tilde{a}).$
As mentioned in Section 2.2, using the log-Sobolev constant $\rho(PP^{*})$ can often significantly improve classical mixing time bounds based on the spectral gap $\lambda(PP^{*})$. Such classical bounds are of the form (Montenegro et al., 2006):
$t_{\operatorname{mix}}^{P}(\epsilon)\leq\frac{C}{\lambda(PP^{*})}\bigg{(}\log\frac{1}{\sqrt{\mu_{\min}}}+\log\frac{1}{\epsilon}\bigg{)},$
where $C$ is some constant. Diaconis and Saloff-Coste (1996) in Lemma $3.1$ show that the log-Sobolev constant $\rho(PP^{*})$ is upper-bounded by the spectral gap $\lambda(PP^{*})$ as follows: $2\rho(PP^{*})\leq\lambda(PP^{*})$. Thus, roughly speaking if
$\log\log\frac{1}{\mu_{\min}}\leq\log\frac{1}{\sqrt{\mu_{\min}}}$ (12)
then the mixing time bound based on the log-Sobolev constant is an improvement upon the mixing time bound based on the spectral gap. To make this more concrete, consider for example a Markov chain on the $d$-dimensional hypercube =$\mathcal{H}=\{-1,1\}^{d}$ with uniform stationary distribution. Then, $\mu_{\min}=2^{-d}$ and clearly Equation (12) is satisfied in this example.
Proof A.2.
of Lemma 2.2 Let the relative entropy be defined as $D(\mu^{t}:\mu):=\sum_{a\in\mathcal{A}^{N}}\mu^{t}(a)\log\frac{\mu^{t}(a)}{\mu(a)}$. Then, for a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ with irreducible transition matrix $P$ the relative entropy $D(\mu^{t}:\mu)$ decays at the following rate (Miclo, 1997):
$D(\mu^{t}:\mu)\leq(1-\rho(PP^{*}))^{t}D(\mu^{0}:\mu).$ (13)
Using Pinsker’s inequality we have that:
$\|\mu^{t}-\mu\|_{TV}\leq\sqrt{\frac{D(\mu^{t}:\mu)}{2}}\leq\sqrt{(1-\rho(PP^{*}))^{t}D(\mu^{0}:\mu)}.$ (14)
Note that $\rho(PP^{*})<1$ since $2\rho(PP^{*})\leq\lambda(PP^{*})$ by Lemma 3.1 in (Diaconis and Saloff-Coste, 1996) and for the spectral gap $\lambda(PP^{*})$ it is known that $\lambda(PP^{*})<1$ (Levin and Peres, 2017). To ensure that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon$, we derive the following lower bound on $t$:
$\sqrt{(1-\rho(PP^{*}))^{t}D(\mu^{0}:\mu)}\leq\epsilon$
[…]
where in line $4$ we used that $\log(1+x)\leq x$ for $x>-1$ and then in line $6$ we used that for any $\mu_{0}$, $D(\mu^{0}:\mu)\leq\log\frac{1}{\mu_{\min}}$ for $\mu_{\min}:=\min_{a\in\mathcal{A}^{N}}\mu(a)$. Thus, we conclude that the mixing time is upper-bounded as follows:
$t_{\operatorname{mix}}^{P}(\epsilon)\leq\frac{1}{\rho(PP^{*})}\bigg{(}\log\log\frac{1}{\mu_{\min}}+2\log\frac{1}{\epsilon}\bigg{)}.$ (15)
Appendix B Convergence of log-linear learning
We first prove Lemma 3.3 and then state a novel lemma which we will use to prove Lemma 3.4. Lastly, we provide a formal proof of the convergence time of log-linear learning in Theorem 3.1.
B.1 Proof of Theorem 3.1
We are now ready to proceed with the proof of Theorem 3.1.
Proof B.1.
of Theorem 3.1 By Lemma 3.4, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
$\rho(PP^{*})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}^{3}}{25N^{2}A^{2}}\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}},$ (16)
where we used that by definition of $P$ in (4) and $\mu$ in (5) $\mu_{\min}$ and $p_{\min}$ can be lower-bounded as follows:
$\mu_{\min}=\min_{a\in\mathcal{A}^{N}}\mu(a)\geq\frac{e^{-\beta}}{A^{N}}$
$P_{a,\tilde{a}}\geq\frac{e^{-\beta}}{NA},\quad\forall\tilde{a}\in\mathcal{A}^{N}(a)\ \Rightarrow\ p_{\min}=\frac{e^{-\beta}}{A}.$
Equation (2) in Lemma 2.2 provides an upper bound on the mixing time:
$t_{\operatorname{mix}}^{P}(\epsilon/4)\leq\frac{1}{\rho(PP^{*})}\bigg{(}\log\log\frac{1}{\mu_{*}}+2\log\frac{4}{\epsilon}\bigg{)}.$
Plugging the bound on the log-Sobolev constant from Equation (16) into the equation above we obtain:
$t_{\operatorname{mix}}^{P}(\epsilon/2)$ $\leq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon}\bigg{)}$
$\leq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log\frac{A^{N}}{e^{-\beta}}+2\log\frac{4}{\epsilon}\bigg{)}$
$\leq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}.$
By setting $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}$ and $\beta$ as in Equation (6) we obtain:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\mathbb{E}_{a\sim\mu}[\Phi(a)]-2\|\mu^{t}-\mu\|_{TV}\max_{a\in\mathcal{A}^{N}}\Phi(a)$
$\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\frac{\epsilon}{2}-\frac{2\epsilon}{4}$
$=\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon,$
where the third line follows from Lemma 3.3, the fact that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/2$, and the fact that $\Phi(\cdot)\in[0,1]$. This concludes the proof of Theorem 3.1.
B.2 Proof of Lemma 3.3
Proof B.2.
[…]
$\mathbb{E}_{a\sim\mu}[\Phi(a)]$ $=\sum_{a\in\mathcal{A}^{N}}\frac{e^{\beta\Phi(a)}}{\sum_{\tilde{a}\in\mathcal{A}^{N}}e^{\beta\Phi(\tilde{a})}}\Phi(a)$
$\geq\sum_{a\in\mathcal{A}^{N}_{*}}\frac{e^{\beta\Phi(a)}}{\sum_{\tilde{a}\in\mathcal{A}^{N}}e^{\beta\Phi(\tilde{a})}}\Phi(a)$
$=\sum_{a\in\mathcal{A}^{N}_{*}}\frac{\Phi(a)}{\sum_{\tilde{a}\in\mathcal{A}^{N}(\epsilon)}e^{\beta(\Phi(\tilde{a})-\Phi(a))}+\sum_{\tilde{a}\in\mathcal{A}^{N}\setminus\mathcal{A}^{N}(\epsilon)}e^{\beta(\Phi(\tilde{a})-\Phi(a))}}$
$\geq\sum_{a\in\mathcal{A}^{N}_{*}}\frac{\Phi(a)}{\sum_{\tilde{a}\in\mathcal{A}^{N}(\epsilon)}e^{0}+\sum_{\tilde{a}\in\mathcal{A}^{N}\setminus\mathcal{A}^{N}(\epsilon)}e^{-\beta\min\{\epsilon,\Delta\}}}$
$\geq\frac{A^{N}_{*}}{A^{N}(\epsilon)+(A^{N}-A^{N}(\epsilon))e^{-\beta\min\{\epsilon,\Delta\}}}\Phi(a^{*}),$ (17)
where the suboptimality gap $\Delta$ is given by $\Delta:=\min_{a\in\mathcal{A}^{N}:\Phi(a)<\Phi(a^{*})}\left(\Phi(a^{*})-\Phi(a)\right)$ with $a^{*}\in\mathcal{A}^{N}_{*}$. Then, we have that:
$\beta\geq\frac{1}{\max\{\epsilon,\Delta\}}$ $\log\bigg{(}(A^{N}-A^{N}(\epsilon))\bigg{(}\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}\bigg{)}\bigg{)}$
[…]
By injecting the last inequality into Equation (17) we deduce that for $\beta\geq\frac{1}{\max\{\epsilon,\Delta\}}\log(A^{N}-A^{N}(\epsilon))(\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}))$, it holds that:
$\mathbb{E}_{a\sim\mu}[\Phi(a)]\geq(1-\epsilon)\max_{a\in A^{N}}\Phi(a)\geq\max_{a\in A^{N}}\Phi(a)-\epsilon,$
where we used that $\Phi(a)\leq 1$ for all $a\in\mathcal{A}^{N}$.
B.3 Proof of Lemma 3.4
We first state the following lemma which will be used to prove Lemma 3.4.
Lemma B.3.
Consider the Markov chain $\{\hat{X}_{t}\}_{t\in\mathbb{N}}$ over a finite state space $\mathcal{A}^{N}$ with transition matrix $\hat{P}\in\mathbb{R}^{A^{N}\times A^{N}}$ given by:
$\hat{P}_{a,\tilde{a}}=\frac{1}{NA}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$ (18)
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. Assuming $A\geq 4$, the log-Soblev constant of $\hat{P}$ is lower-bounded by:
$\rho(\hat{P})\geq\frac{4\pi^{2}}{25NA}.$
Proof B.4.
We will lower bound the log-Soblev constant of $\hat{P}$ in terms of the log-Soblev constant of another Markov chain for which a lower bound on the log-Soblev constant is known.
First, note that the Markov chain $\{\hat{X}_{t}\}_{t\in\mathbb{N}}$ is aperiodic and irreducible with stationary distribution $\hat{\mu}(a)=1/A^{N}$. This can be verified by checking the detailed balance equations given by $\hat{\mu}(a)\hat{P}_{a,\tilde{a}}=\hat{\mu}(\tilde{a})\hat{P}_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$. Next, we make use of Corollary $2.15$ in (Montenegro et al., 2006) to lower-bound $\rho(\hat{P})$ in terms of the log-Soblev constant of another Markov chain $\bar{X}_{t}$ with transition matrix $\bar{P}$ and stationary distribution $\bar{\mu}$. Namely,
$\rho(\hat{P})\geq\frac{1}{MC}\rho(\bar{P}),$
where
$M=\max_{a\in\mathcal{A}^{N}}\frac{\hat{\mu}(a)}{\bar{\mu}(a)},$
$C=\max_{a\neq\tilde{a}:\hat{P}_{a,\tilde{a}}\neq 0}\frac{\bar{\mu}(a)\bar{P}_{a,\tilde{a}}}{\hat{\mu}(a)\hat{P}_{a,\tilde{a}}}.$
As the comparison Markov chain, we consider the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ with $\bar{X}_{t}=\prod_{i=1}^{N}\bar{X}_{i,t}$ on the state space $\mathbb{Z}_{K^{N}}=\prod_{i=1}^{N}\mathbb{Z}_{K}$. Here, each $\{\bar{X}_{i_{t}}\}_{t\in\mathbb{N}}$ is a simple random walk on the finite circle $\mathbb{Z}_{K}=\{1,\ldots,K\}$ with $K\geq 4$. The transition matrix and the stationary distribution of the simple random walk $\bar{X}_{i_{t}}$ are given by $\bar{P}_{i_{k,k\pm 1}}=\frac{1}{2}$ and $\bar{\mu}_{i}(k)=1/K$, respectively. Furthermore, the log-Soblev constant $\rho(\bar{P}_{i})$ is lower bounded by $\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25K^{2}}$ (Diaconis and Saloff-Coste, 1996, Example 4.2). Thus, by definition the product chain $\bar{X}_{t}$ (Diaconis and Saloff-Coste, 1996, Sec. 2.5) has the following transition matrix:
$\bar{P}_{\textbf{k},\tilde{\textbf{k}}}=\frac{1}{2N}\mathds{1}_{\tilde{\textbf{k}}=(k_{i}\pm 1,\textbf{k}_{-i})},$
and stationary distribution:
$\bar{\mu}(\textbf{k})=\prod_{i=1}^{N}\bar{\mu}_{i}(k_{i})=\prod_{i=1}^{N}\frac{1}{K}=\frac{1}{K^{N}}.$
Furthermore, by Lemma $3.2$ in (Diaconis and Saloff-Coste, 1996) the log-Soblev constant $\rho(\bar{P})$ of the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ is lower bounded by:
$\rho(\bar{P})=\frac{1}{N}\min_{i\in\{1,\ldots,N\}}\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25NK^{2}}.$ (19)
Next, note that there is a one-to-one mapping between the set $A^{N}$ and the set $\mathbb{Z}_{K}$ with $|\mathcal{A}|=A=K$ and thus a one-to-one mapping between the set $\mathcal{A}^{N}$ and the set $\mathbb{Z}_{K^{N}}$ with with $|\mathcal{A}^{N}|=A^{N}=K^{N}$. Therefore, we can assume that the Markov chains $\hat{X}_{t}$ and $\bar{X}_{t}$ operate on the same state space. By Equation (19) and with the following upper bounds on $M$ and $C$:
$M=\max_{a\in\mathcal{A}^{N}}\frac{\hat{\mu}(a)}{\bar{\mu}(a)}=\frac{A^{N}}{A^{N}}=1$
$C=\max_{a\neq\tilde{a}:\hat{P}_{a,\tilde{a}}\neq 0}\frac{\bar{\mu}(a)\bar{P}_{a,\tilde{a}}}{\hat{\mu}(a)\hat{P}_{a,\tilde{a}}}=\frac{A}{2}.$
the log-Soblev constant $\rho(\hat{P})$ can be lower-bounded by:
$\rho(\hat{P})\geq\frac{1}{MC}\rho(\bar{P})\geq\frac{16\pi^{2}}{25NA^{3}},$
which concludes the proof.
Next, we proceed to prove Lemma 3.4.
Proof B.5.
of Lemma 3.4 The Markov chain $X_{t}$ with transition matrix $P$ defined in (7) is aperiodic and irreducible and thus a unique stationary distribution $\mu$ exists with $\mu^{t}=\mu^{0}P^{t}\rightarrow\mu$ for $t\rightarrow\infty$ from any initial distribution $\mu^{0}$. We define $\mu_{\max}$ and $\mu_{\min}$ as $\max_{a\in\mathcal{A}}\mu(a)\leq\mu_{\max}\leq 1$ and $\min_{a\in\mathcal{A}}\mu(a)\geq\mu_{\min}>0$, where $\mu_{\min}>0$ follows from the irreducibility of $X_{t}$.
Next, we consider the modified Markov chain $\{X^{*}_{t}\}_{t\in\mathbb{N}}$ with transition matrix $PP^{*}$ which is aperiodic and irreducible since $X_{t}$ is aperiodic and irreducible. Concretely, since $P$ contains self-loops, i.e., $P_{a,a}>0$, it follows that $PP^{*}$ contains self-loops, i.e.,
$PP^{*}_{a,a}=\sum_{a^{\prime}\in\mathcal{A}}P_{a,a^{\prime}}P^{*}_{a^{\prime},a}=\sum_{a^{\prime}\in\mathcal{A}}P_{a,a^{\prime}}\frac{\mu(a)P_{a,a^{\prime}}}{\mu(a^{\prime})}\geq P_{a,a}P_{a,a}>0,$
and thus $X^{*}_{t}$ is aperiodic. Furthermore, for any $a,\tilde{a}\in\mathcal{A}^{N}$:
$(PP^{*})^{N}_{a,\tilde{a}}$ $=\sum_{\begin{subarray}{c}a_{l}\in\mathcal{A}^{N}\\ l=1,\ldots,N-1\end{subarray}}(PP^{*})_{a,a_{1}}\ldots(PP^{*})_{a_{N-1},\tilde{a}}$
$=\sum_{\begin{subarray}{c}a_{l}\in\mathcal{A}^{N}\\ l=1,\ldots,N-1\end{subarray}}\sum_{a^{\prime}\in\mathcal{A}^{N}}P_{a,a^{\prime}}P^{*}_{a^{\prime},a_{1}}\ldots\sum_{a^{\prime}\in\mathcal{A}^{N}}P_{a_{N-1},a^{\prime}}P^{*}_{a^{\prime},\tilde{a}}$
$\geq\sum_{\begin{subarray}{c}a_{l}\in\mathcal{A}^{N}\\ l=1,\ldots,N-1\end{subarray}}P_{a,a_{1}}P_{a_{1},a_{1}}\ldots P_{a_{N-1},\tilde{a}}P_{\tilde{a},\tilde{a}}>0,$
where we used that $P^{N}_{a,\tilde{a}}>0$ and $P_{a,a}>0$ for all $a,\tilde{a}\in\mathcal{A}^{N}$ as well as the identity $\mu(a)P^{*}_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$. It follows that $X^{*}_{t}$ is irreducible. Thus, for $X^{*}_{t}$ a unique stationary distribution $\mu^{*}$ exists. By Proposition $1.23$ in (Levin and Peres, 2017) the stationary distribution of $P^{*}$ is given by $\mu$ and thus the stationary distribution of $PP^{*}$ is given by $\mu$ since $\mu PP^{*}=\mu P^{*}=\mu$. Furthermore, the following holds for the transition matrix $PP^{*}$ :
$\frac{1}{N}p_{\min}^{2}\mathds{1}_{\tilde{a}\in\mathcal{A}^{N}(a)}\leq(PP^{*})_{a,\tilde{a}}\leq\mathds{1}_{\tilde{a}\in\mathcal{A}^{N}(a)}$
where
$(PP^{*})_{a,\tilde{a}}=\sum_{a^{\prime}\in\mathcal{A}}P_{a,a^{\prime}}P^{*}_{a^{\prime},\tilde{a}}\geq P_{a,\tilde{a}}P^{*}_{\tilde{a},\tilde{a}}\geq P_{a,\tilde{a}}P^{*}_{\tilde{a},\tilde{a}}\geq\frac{p_{\min}^{2}}{N}.$
Next, we compute $M$ and $C$, defined bellow, of the Markov chains $X^{*}_{t}$ and $X_{t}$:
$M=\max_{a\in\mathcal{A}^{N}}\frac{\mu(a)}{\mu(a)}=1$
$C=\max_{a\neq\tilde{a}:(PP^{*})_{a,\tilde{a}}\neq 0}\frac{\mu(a)P_{a,\tilde{a}}}{\mu(a)(PP^{*})_{a,\tilde{a}}}\leq\frac{N}{p_{\min}^{2}}.$
From Corollary $2.15$ in (Diaconis and Saloff-Coste, 1996), it follows that the log-Soblev constant $\rho(PP^{*})$ is lower-bounded by:
$\rho(PP^{*})\geq\frac{1}{MC}\rho(P)\geq\frac{p_{\min}^{2}}{N}\rho(P).$ (20)
Next, we compare the Markov chain $X_{t}$ to the Markov chain $\hat{X}_{t}$ with transition matrix $\hat{P}$ specified in Equation (18) of Lemma B.3. To this end, we compute $M$ and $C$ of the Markov chains $X_{t}$ and $\hat{X}_{t}$:
$M=\max_{a\in\mathcal{A}^{N}}\frac{\mu(a)}{\hat{\mu}(a)}\leq A^{N}$
$C=\max_{a\neq\tilde{a}:P_{a,\tilde{a}}\neq 0}\frac{\hat{\mu}(a)\hat{P}_{a,\tilde{a}}}{\mu(a)P_{a,\tilde{a}}}\leq\frac{N}{A^{N}NA\mu_{\min}p_{\min}},$
Thus, by Corollary $2.15$ in (Diaconis and Saloff-Coste, 1996) and by Lemma B.3 the log-Soblev constant $\rho(P)$ can be lower-bounded by:
$\rho(P)\geq\frac{1}{MC}\rho(\hat{P})\geq A^{N}A\mu_{\min}p_{\min}\rho(\hat{P})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}}{25NA^{2}}.$ (21)
Combining Equation (20) and (21), we conclude that the log-Soblev constant $\rho(PP^{*})$ is lower-bounded by:
$\rho(PP^{*})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}^{3}}{25N^{2}A^{2}}.$
B.4 Modified log-linear learning in symmetric potential games
In the following, we consider a symmetric potential game and show that in this setting, we obtain convergence time guarantees to an $\epsilon$-efficient NE that depend polynomially on the number of players $N$ and on $1/\epsilon$. This improves the convergence time result provided in (Shah and Shin, 2010) which shows an exponential dependence on $1/\epsilon$.
A game is said to be symmetric if it satisfies the following definition.
Definition B.6.
A game is symmetric if for any permutation $\pi$ of $\{1,\ldots,N\}$ the following holds:
$U_{i}(a_{1},\ldots,a_{N})=U_{\pi(i)}(a_{\pi(1)},\ldots,a_{\pi(N)}).$
In other words, the utility of player $i$ depends only on how many players are playing each action $a\in\mathcal{A}$ and not on players’ identities. Thus, in a symmetric potential game, if $A<N$, the potential function $\Phi$ can be redefined in terms of a lower-dimensional function $\Phi:\Psi^{\mathcal{A}}\rightarrow[0,1]$, where
$\Psi^{\mathcal{A}}:=\bigg{\{}\bigg{(}\frac{v_{1}}{N},\ldots,\frac{v_{A}}{N}\bigg{)}\>|\>v_{j}\in\mathbb{Z}_{+}\>\forall j\in[A],\>\sum_{j=1}^{A}v_{j}=N\bigg{\}}$ (22)
with cardinality $Y=|\Psi^{\mathcal{A}}|\leq(N+1)^{A-1}$. Then, for any $a=(a_{1},\ldots,a_{N})\in\mathcal{A}^{N}$ we have $\Phi(a)=\Phi(x(a))$ with $x(a)=(x_{1}(a),\ldots,x_{N}(a))$. Here $x_{j}(a)$ denotes the fraction of players that selected action $j\in A^{N}$, i.e., $x_{j}(a)=1/N|\{i\in[N]\>|\>a_{i}=j\}|$.
To obtain their results for symmetric potential games, Shah and Shin (2010) propose modified log-linear learning, a variant of log-linear learning. In modified log-linear learning every player $i$ has an independent exponential clock of rate $\alpha/z_{i}^{t}$, where $\alpha>0$ is a parameter and $z_{i}^{t}:=1/N|\{j\in[N]\>|\>a_{j}^{t}=a_{i}^{t}\}|$. [^8] The only change compared to log-linear learning is that in log-linear learning players have a fixed exponential clock rate of $1$. This means that the times between two consecutive clock-ticks are independent and distributed as the exponential distribution of mean $\alpha/z_{i}^{t}$. When the clock of player $i$ ticks, she is allowed to alter her current action. Player $i$ samples an action from her strategy $p_{i}^{t}\in\Delta(\mathcal{A})$ defined as in Equation (3). Modified log-linear learning induces an aperiodic and irreducible Markov chain on the lower-dimensional state space $\Psi^{\mathcal{A}}$ with stationary distribution $\mu_{m}\in\Delta(\Psi^{\mathcal{A}})$ given by (Shah and Shin, 2010, Lemma 2):
$\mu_{m}(x)=\frac{e^{\beta\Phi(x)}}{\sum_{\tilde{x}\in\Psi_{N}^{\mathcal{A}}}e^{\beta\Phi(\tilde{x})}}\quad\forall x\in\Psi^{\mathcal{A}}.$
In the following, we show that in a symmetric potential game, if all players adhere to modified log-linear learning an $\epsilon$-efficient NE is reached in time polynomial in $N$ and $1/\epsilon$.
Corollary B.7.
Consider a symmetric potential game with potential function $\Phi:\Psi^{A^{N}}\rightarrow[0,1]$. For any $\epsilon\in(0,1)$ and any initial distribution $\mu_{m}^{0}$, assume that players adhere to modified log-linear learning with:
$\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\left(\frac{N^{A}}{\epsilon}\right)\right).$
Then,
$\mathbb{E}_{x\sim\mu_{m}^{t}}[\Phi(x)]\geq\max_{x\in\Psi^{\mathcal{A}}}\Phi(x)-\epsilon,$
for
$t\geq\frac{N}{\alpha c}e^{3\beta}\left(\log((A-1)\log(N+1))+\log\beta+2\log\frac{4}{\epsilon}\right)=\tilde{\mathcal{O}}\left(N\left(\frac{N^{A}}{\epsilon}\right)^{\frac{1}{\max\{\epsilon,\Delta\}}}\right),$
where $c$ is a constant that depends on $A$.
Proof B.8.
[…]
$\mathbb{E}_{x\sim\mu_{m}}[\Phi(x)]$ $=\sum_{a\in\Psi^{\mathcal{A}}}\frac{e^{\beta\Phi(x)}}{\sum_{\tilde{x}\in\Psi^{\mathcal{A}}}e^{\beta\Phi(\tilde{x})}}\Phi(x)$
$=\sum_{a\in\Psi_{*}^{\mathcal{A}}}\frac{1}{\sum_{\tilde{x}\in\Psi^{\mathcal{A}}(\epsilon)}e^{\beta(\Phi(\tilde{x})-\Phi(x))}+\sum_{\tilde{x}\in\Psi^{\mathcal{A}}\setminus\Psi^{\mathcal{A}}(\epsilon)}e^{\beta(\Phi(\tilde{x})-\Phi(x))}}\Phi(x)$
$\geq\sum_{a\in\Psi_{*}^{\mathcal{A}}}\frac{1}{\sum_{\tilde{x}\in\Psi^{\mathcal{A}}(\epsilon)}e^{0}+\sum_{\tilde{x}\in\Psi^{\mathcal{A}}\setminus\Psi^{\mathcal{A}}(\epsilon)}e^{-\beta\min\{\epsilon,\Delta\}}}\Phi(x)$
$\geq\frac{Y_{*}}{Y(\epsilon)+(Y-Y(\epsilon))e^{-\beta\min\{\epsilon,\Delta\}}}\Phi(x^{*}),$
where the suboptimality gap $\Delta$ is given by $\Delta:=\min_{x\in\Psi^{\mathcal{A}}:\Phi(x)<\Phi(x^{*})}\left(\Phi(x^{*})-\Phi(x)\right)$ with $x^{*}\in\Psi_{*}^{\mathcal{A}}$. Then, for
$\beta\geq$ $\frac{1}{\max\{\epsilon,\Delta\}}\log\bigg{(}(N+1)^{A-1}\bigg{(}\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}\bigg{)}\bigg{)},$ (23)
it holds that
$\frac{Y_{*}}{Y(\epsilon)+(Y-Y(\epsilon))e^{-\beta\min\{\epsilon,\Delta\}}}\geq 1-\epsilon,$
where we used that $Y\leq(N+1)^{A-1}$. We deduce that for $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log(\frac{N^{A}}{\epsilon}))$, it holds that:
$\mathbb{E}_{x\sim\mu_{m}}[\Phi(x)]\geq(1-\epsilon)\max_{x\in\Psi^{\mathcal{A}}}\Phi(x)\geq\max_{x\in\Psi^{\mathcal{A}}}\Phi(x)-\epsilon.$
The proof now follows from the same analysis as in the proof of Theorem $3$ in (Shah and Shin, 2010) with the exception that we do not make use of Lemma $6$ in (Shah and Shin, 2010) but replace it with our analysis above. Concretely, we set $\beta$ as specified in Equation (23) rather than as in (Shah and Shin, 2010, Eq. (8)).
Appendix C Binary log-linear learning
Binary log-linear learning induces an irreducible and aperiodic Markov chain $\{X_{t}\}_{t\in\mathbb{Z}_{+}}$ with a time-reversible transition matrix $P\in\mathbb{R}^{\mathcal{A}\times\mathcal{A}}$ given by:
$P_{a,\tilde{a}}=\frac{1}{N}\frac{1}{A}\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{e^{\beta U_{i}(a_{i},\tilde{a}_{-i})}+e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$ (24)
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The additional term $1/A$ stems from the fact that player $i$ first randomly samples an action $\tilde{a}_{i}$ and then decides between this action and her previous action. Arslan et al. (2007) show that its stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ is given by :
$\mu(a)=\frac{e^{\beta\Phi(a)}}{\sum_{\tilde{a}\in A^{N}}e^{\beta\Phi(\tilde{a})}}\quad\forall a\in\mathcal{A}^{N}.$ (25)
The stationary distribution of binary log-linear learning is the same as that of log-linear learning (Equation (5)). Thus, log-linear- and binary log-linear learning converge to an approximately efficient Nash equilibrium in the long run. We briefly outline the proof of Theorem 4.1 and then provide a detailed proof.
Proof C.1.
The proof follows from the same line of arguments as the proof of Theorem 3.1. In particular, the first step in the proof of Theorem 3.1 remains the same since binary log-linear learning has the same stationary distribution as log-linear learning. In the second step in the proof of Theorem 3.1, the main difference is that the transition matrix in (24) of binary log-linear learning differs from the transition matrix in (4) of log-linear learning. Thus, the log-Sobolev constant of binary log-linear can be lower-bounded as follows:
$\rho(PP^{*})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}^{3}}{25N^{2}A^{2}}\geq\frac{2\pi^{2}e^{-4\beta}}{25N^{2}A^{5}},$ (26)
while the log-Sobolev constant of log-linear can be lower-bounded as follows:
$\rho(PP^{*})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}^{3}}{25N^{2}A^{2}}\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}}.$
Then, we use Lemma 2.2 to show that
$\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4$
for $t\geq\frac{1}{\rho(PP^{*})}(\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon})$ with $\rho(PP^{*})$ lower-bounded as in Equation (26).
Proof C.2.
of Theorem 4.1 By Lemma 3.4, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
$\rho(PP^{*})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}^{3}}{25N^{2}A^{2}}\geq\frac{2\pi^{2}e^{-4\beta}}{25N^{2}A^{5}},$
where we used that by definition of $P$ in (24) and $\mu$ in (25) $\mu_{\min}$ and $p_{\min}$ can be lower-bounded as follows:
$\mu_{\min}=\min_{a\in\mathcal{A}^{N}}\mu(a)\geq\frac{e^{-\beta}}{A^{N}}$
$P_{a,\tilde{a}}\geq\frac{e^{-\beta}}{N2A},\quad\forall\tilde{a}\in\mathcal{A}^{N}(a)\ \Rightarrow\ p_{\min}=\frac{e^{-\beta}}{2A}.$
Equation (2) in Lemma 2.2 provides the following upper bound on the mixing time:
$t_{\operatorname{mix}}^{P}(\epsilon/4)\leq\frac{1}{\rho(PP^{*})}\bigg{(}\log\log\frac{1}{\mu_{*}}+2\log\frac{4}{\epsilon}\bigg{)}.$
Plugging the bound on the log-Sobolev constant into this equation we obtain:
$t_{\operatorname{mix}}^{P}(\epsilon/4)$ $\leq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\bigg{(}\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon}\bigg{)}$
$\leq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\bigg{(}\log\log\frac{A^{N}}{e^{-\beta}}+2\log\frac{4}{\epsilon}\bigg{)}$
$\leq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}.$
Set $t$ as:
$t\geq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}$ (27)
and set $\beta$ as:
$\beta$ $\geq\frac{1}{\max\{\epsilon/2,\Delta\}}\log\bigg{(}(A^{N}-A^{N}(\epsilon/2))\bigg{(}\frac{4}{\epsilon A^{N}(\epsilon/2)}-\frac{1}{A^{N}(\epsilon/2)}\bigg{)}\bigg{)}.$ (28)
Then, we obtain the following upper bound:
$\mathbb{E}[\Phi(a^{t})]$ $=\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$
$\geq\mathbb{E}_{a\sim\mu}[\Phi(a)]-2\|\mu^{t}-\mu\|_{TV}\max_{a\in\mathcal{A}^{N}}\Phi(a)$
$\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\frac{\epsilon}{2}-\frac{2\epsilon}{4}$
$=\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon,$
where the third line follows from Lemma 3.3, the fact that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4$ for $t$ set as in Equation (27), and the fact that $\Phi(\cdot)\in[0,1]$. Lemma 3.3 is applicable when all players adhere to binary-based log-linear learning rather than log-linear learning since the proof of Lemma 3.3 depends only on the stationary distribution $\mu$ of the corresponding learning rule which is the same for log-linear learning and binary log-linear learning. This concludes the proof of Theorem 4.1.
Appendix D Robustness of log-linear learning
In this section, we prove Theorem 4.3 and apply it to the corrupted-utility case to prove Corollary 4.4.
D.1 Proof of Theorem 4.3
Proof D.1.
Consider a learning rule $P$, to prove Theorem 4.3, we first provide a decomposition that relates the expected value of the potential when the agents follow $P$ defined in Equation (8) to the same quantity when the agents follow $P_{\ell}$ defined in Equation (4) instead.
We have for all $t,t^{\prime}\in\mathbb{N}$ that:
$\mathbb{E}_{\mu_{0}P^{t}}[\Phi]$ $=\mathbb{E}_{\mu_{0}P_{\ell}^{t^{\prime}}}[\Phi]+\mathbb{E}_{\mu_{0}P^{t}}[\Phi]-\mathbb{E}_{\mu_{0}P_{\ell}^{t^{\prime}}}[\Phi]$
$\geq\mathbb{E}_{\mu_{0}P_{\ell}^{t^{\prime}}}[\Phi]-\sqrt{A^{N}}\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ (29)
where the last line follows because $|\Phi(a)|\leq 1$ for all $a\in\mathcal{A}^{N}$ and because .$\|.\|_{1}\leq\sqrt{A^{N}}\|\cdot\|_{2}$. The rest of the proof is based on controlling $\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ using our mixing time results.
Decomposition: Using Lemma 4.2 we obtain:
$\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ $\leq\|P^{t}-\mu\|_{2}+\|P_{\ell}^{t^{\prime}}-\mu_{\ell}\|_{2}+\|\mu-\mu_{\ell}\|_{2}$
$\leq\|P^{t}-\mu\|_{2}+\|P_{\ell}^{t^{\prime}}-\mu_{\ell}\|_{2}+L(P_{\ell})\|P-P_{\ell}\|_{2}$
$\leq 2\|P^{t}-\mu\|_{TV}+\|P_{\ell}^{t^{\prime}}-\mu_{\ell}\|_{2}+L(P_{\ell})\|P-P_{\ell}\|_{2},$
where $L(P_{\ell})=\frac{2A^{N}}{\rho(P_{\ell}P_{\ell}^{*})}(\log\log\frac{1}{\mu_{\ell,\min}}+\log(8A^{N}))$ follows from Lemma 4.2. In Theorem 3.1, we showed that $\mu_{\ell,\min}\geq\frac{e^{-\beta}}{A^{N}}$ and $\rho(P_{\ell}P_{\ell}^{*})\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}}$, therefore
$L(P_{\ell})\leq\frac{25N^{2}A^{N+5}e^{4\beta}}{8\pi^{2}}(\log\log A^{N}e^{\beta}+\log(8A^{N})).$
Thus, for
$\begin{cases}t&\geq t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}})))\\ t^{\prime}&\to\infty\\ \beta&=\log\bigg{(}(A^{N}-A^{N}(\epsilon/2))\bigg{(}\frac{4}{\epsilon/2A^{N}(\epsilon/2)}-\frac{1}{A^{N}(\epsilon/2)}\bigg{)}\bigg{)}/\max\{\epsilon/2,\Delta\}\end{cases}$
we have
$\begin{cases}\|P^{t}-P_{\ell}^{t}\|_{2}&\leq\epsilon/\left(2\sqrt{A^{N}}\right)+L(P_{\ell})\|P-P_{\ell}\|_{2}\\ \mathbb{E}_{a\sim\mu^{0}P_{\ell}^{t^{\prime}}}[\Phi(a)]&\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon/2,\\ L(P_{\ell})&=\mathcal{O}\Bigg{(}N^{2}A^{N+5}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}\Bigg{(}\log\log A^{N}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}+\log(A^{N})\Bigg{)}\Bigg{)},\end{cases}$
where the second line follows from Lemma 3.3. Plugging the above inequalities with the bound on $L$ from Lemma 4.2 into the decomposition (29) proves the desired result. We now bound the mixing time $t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}}))$ of the Markov chain induced by $P$.
Mixing time bound: To bound the mixing time of the Markov chain induced by $P$, we use inequality (2) and Lemma 3.4. Assuming a lower bound of $p_{\min}/N$ on the probabilities of all feasible transitions implies a bound on the stationary distribution as we will show next.
Lower bound $(\mu_{P})_{\min}$: Since $P$ has a positive probability of transitioning from $a\in\mathcal{A}^{N}$ to any $\tilde{a}\in\mathcal{N}(a)$, it follows that the corresponding $N$-step transition $P^{N}$ has a positive probability of transitioning from any $a\in\mathcal{A}^{N}$ to any $a^{\prime}\in\mathcal{A}^{N}$, i.e.,
$\forall a,a^{\prime}\in\mathcal{A}^{N}:\quad P^{N}_{a,a^{\prime}}\geq N!\>(p_{\min}/N)^{N}.$
Note that the least probable transitions are when $a$ and $a^{\prime}$ are such that $\forall i\in[N]:\>a_{i}\neq a^{\prime}_{i}$. For all such transitions, the possible paths using $P^{N}$ are the permutations of $\{1,\ldots,N\}$ (each of the $N$ steps is a new player updating their action). There are $N!$ such permutations and each player $i\in[N]$ can update $a_{i}$ to $a^{\prime}_{i}$ with probability larger than $p_{\min}/N$.
Since $P$ is an irreducible and aperiodic transition matrix, the Markov chain induced by $P$ has a unique stationary distribution $\mu_{P}$. It is known that the Markov chain induced by $P^{N}$ has the same stationary distribution as the one induced by $P$. Therefore, we have for all $a\in\mathcal{A}^{N}$:
$\mu_{P}(a)$ $=\sum_{\tilde{a}\in\mathcal{A}^{N}}P^{N}_{\tilde{a},a}\mu_{P}(\tilde{a})$
$\geq\sum_{\tilde{a}\in\mathcal{A}^{N}}N!\>(p_{\min}/N)^{N}\mu_{P}(\tilde{a})=N!\>(p_{\min}/N)^{N}.$
Therefore, $(\mu_{P})_{\min}\geq N!\>(p_{\min}/N)^{N}$.
Deducing the mixing-time bound: We can now give an explicit bound on the mixing time of the transition $P$. First, we have by Lemma 3.4:
$\rho(PP^{*})$ $\geq\frac{16\pi^{2}A^{N}(\mu_{P})_{\min}p_{\min}^{3}}{25N^{2}}$
$\geq\frac{4\pi^{2}A^{N}p_{\min}^{N+3}N!}{25N^{N+2}}.$
Then, using Stirling’s formula, we have $N!\geq\sqrt{2\pi N}\left(\frac{N}{e}\right)^{N}$, therefore
$\rho(PP^{*})\geq\frac{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}{25N^{3/2}e^{N}}$
Then, using inequality (2), we have:
$t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}}))$ $\leq\frac{1}{\rho(PP^{*})}\Bigg{(}\log\log\frac{1}{(\mu_{P})_{\min}}+2\log\frac{4\sqrt{A^{N}}}{\epsilon}\Bigg{)}$
$\leq\frac{25N^{3/2}e^{N}}{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}\Bigg{(}\log\log\frac{e^{N}}{p_{\min}^{N}\sqrt{2\pi N}}+2\log\frac{4\sqrt{A^{N}}}{\epsilon}\Bigg{)}.$
This concludes the proof of Theorem 4.3.
D.2 Proof of Corollary 4.4
The key idea is to show that the transition matrix of the Markov chain induced by corrupted utilities is close to its corruption-free counterpart.
Proof D.2.
If all players adhere to log-linear learning with corrupted utilities, the induced Markov chain’s transition matrix $\hat{P}$ is given, for all $a,\tilde{a}\in\mathcal{A}^{N}$ by:
$\hat{P}_{a,\tilde{a}}$ $=\frac{1}{N}\frac{e^{\beta\hat{U}_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}_{i}}e^{\beta\hat{U}_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)},$
$=\frac{1}{N}\frac{e^{\beta(U_{i}(\tilde{a}_{i},\tilde{a}_{-i})+\xi_{i}(\tilde{a}_{i},\tilde{a}_{-i}))}}{\sum_{a_{i}^{\prime}\in\mathcal{A}_{i}}e^{\beta(U_{i}(a_{i}^{\prime},\tilde{a}_{-i})+\xi_{i}(a_{i}^{\prime},\tilde{a}_{-i}))}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}.$
Since we assumed that the noise is bounded, we can deduce that
$P_{a,\tilde{a}}e^{-2\beta\xi}\leq P_{a,\tilde{a}}\leq P_{a,\tilde{a}}e^{2\beta\xi},$
where $P_{a,\tilde{a}}=\frac{1}{N}\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}_{i}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$ is the transition with the noise-free utility. This entails that
$P_{a,\tilde{a}}(e^{-2\beta\xi}-1)\leq\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\leq P_{a,\tilde{a}}(e^{2\beta\xi}-1),$
then, since $e^{-2\beta\xi}-1<0$ and $P_{a,\tilde{a}}\leq 1/N$ for all $a,\tilde{a}\in\mathcal{A}^{N}$, we deduce that
$(e^{-2\beta\xi}-1)/N\leq\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\leq(e^{2\beta\xi}-1)/N,$
and
$|\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}|\leq\frac{1}{N}\max\left\{e^{2\beta\xi}-1,1-e^{-2\beta\xi}\right\},$
Finally, since $2\beta\xi\leq 1$ and by using that: $01-e^{-x}<x\text{ for }x>$ 0, and that: $e^{x}-1<\frac{7}{4}x\text{ for }x\in[0,1]$. Then,
$|\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}|\leq\frac{1}{N}\max\left\{\frac{7}{2}\beta\xi,2\beta\xi\right\}=\frac{7}{2N}\beta\xi,$
and finally
$\|\hat{P}-P\|_{2}$ $\leq\sqrt{\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\frac{49}{4N^{2}}\beta^{2}\xi^{2}}$
$=\frac{7A^{N}}{2N}\beta\xi.$
Also, since $P_{a,\tilde{a}}\geq P_{a,\tilde{a}}e^{-2\beta\xi}$ and using $P_{a,\tilde{a}}\geq\frac{e^{-\beta}}{NA}$ then we deduce that $P_{a,\tilde{a}}\geq\frac{e^{-\beta(1+2\xi)}}{NA}$. We conclude the proof with a straightforward application of Theorem 4.3 with $p_{\min}=e^{-\beta(1+2\xi)}/A$ and $\|\hat{P}-P\|_{2}\leq\frac{7A^{N}}{2N}\beta\xi$.
D.3 Proof of Corollary 4.5
Similar to Corollary 4.4, we proceed by showing that the transition matrix of the Markov chain induced by fixed-share log-linear learning is close to that of log-linear learning.
Proof D.3.
If all players adhere to fixed-share log-linear learning, the induced Markov chain’s transition matrix $\hat{P}$ is given, for all $a,\tilde{a}\in\mathcal{A}^{N}$ by:
$\hat{P}_{a,\tilde{a}}=\frac{1}{N}\left(\frac{\xi}{A}+\frac{(1-\xi)e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}.$ (30)
Then, we have that
$\hat{P}_{a,\tilde{a}}\geq\left(\frac{\xi}{NA}+\frac{(1-\xi)e^{-\beta}}{NA}\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)},$
which entails that $\hat{P}$ satisfies the condition of Theorem 4.3 with $p_{\min}\geq\frac{\xi}{A}+\frac{(1-\xi)e^{-\beta}}{A}$. Additionally, we can show that:
$\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}$ $=\frac{1}{N}\left(\frac{\xi}{A}-\frac{\xi e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$
$=\frac{\xi}{N}\left(\frac{1}{A}-\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)},$
where $P$ is the transition matrix of log-linear learning. Therefore,
$\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\right)^{2}$ $=\frac{\xi^{2}}{N^{2}}\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\frac{1}{A^{2}}-\frac{2e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{A\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}+\frac{e^{2\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\left(\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}\right)^{2}}\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$
$\leq\frac{\xi^{2}}{N^{2}}\sum_{a\in\mathcal{A}^{N}}\left(\frac{N}{A}-\frac{2N}{A}+N\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$
$\leq NA^{N},$
[…]
Finally, since the spectral norm is smaller than the Frobenius norm, then
$\|\hat{P}-P\|_{2}$ $\leq\sqrt{\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\right)^{2}}$
$\leq\xi\sqrt{\frac{A^{N}}{N}}.$
The proof is then concluded by a straightforward application of Theorem 4.3 with $p_{\min}\geq\frac{\xi}{A}+\frac{(1-\xi)e^{-\beta}}{A}$ and $\|\hat{P}-P\|_{2}\leq\xi\sqrt{\frac{A^{N}}{N}}$.
D.4 Proof of Lemma 4.2
Proof D.4.
Denote by $M\in\mathbb{R}^{A^{N}\times A^{N}}$ the matrix, where each row corresponds to $\mu_{1}$. For all $t\in\mathbb{N}$, we have that:
$\mu_{1}-\mu_{2}$ $=(P_{1}^{t})^{\top}(\mu_{1}-\mu_{2})+(P_{1}^{t}-P_{2}^{t})^{\top}\mu_{2}$
$=(P_{1}^{t}-M)^{\top}(\mu_{1}-\mu_{2})+M^{\top}(\mu_{1}-\mu_{2})$
$\quad+(P_{1}^{t}-P_{2}^{t})^{\top}\mu_{2}.$
This yields:
$\|\mu_{1}-\mu_{2}\|_{2}$ $\leq\|(P_{1}^{t}-M)^{\top}(\mu_{1}-\mu_{2})\|_{2}$
$\quad+\|M^{\top}(\mu_{1}-\mu_{2})\|_{2}+\|(P_{1}^{t}-P_{2}^{t})^{\top}\|_{2}\|\mu_{2}\|_{2}$
$\leq\|P_{1}^{t}-M\|_{2}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+\|M^{\top}(\mu_{1}-\mu_{2})\|_{2}+\|P_{1}^{t}-P_{2}^{t}\|_{2}$
$\leq 2\sqrt{A^{N}}\|P_{1}^{t}-M\|_{TV}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+\|M^{\top}(\mu_{1}-\mu_{2})\|_{2}+\|P_{1}^{t}-P_{2}^{t}\|_{2}$
where in the last inequality we used the equivalence of $\|\cdot\|_{2}$ and $\|\cdot\|_{1}$ and that $\|\cdot\|_{1}=2\|\cdot\|_{TV}$ by definition of the total variation distance. Furthermore:
$M^{\top}(\mu_{1}-\mu_{2})$ $=\bigg{(}\mu_{1}(a)\underbrace{\sum_{a^{\prime}\in\mathcal{A}^{N}}(\mu_{1}(a^{\prime})-\mu_{2}(a^{\prime}))}_{=0}\bigg{)}_{a\in\mathcal{A}^{N}}$
$=0.$
Therefore, we obtain that:
$\|\mu_{1}-\mu_{2}\|_{2}$ $\leq 2\sqrt{A^{N}}\|P_{1}^{t}-M\|_{TV}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+\|P_{1}^{t}-P_{2}^{t}\|_{2}.$ (31)
Note that for the second term on the right-hand side in the equation above we have:
$P_{1}^{t}-P_{2}^{t}$ $=P_{1}^{t}+\sum_{l=1}^{t-1}(P_{1}^{t-l}P_{2}^{l}-P_{1}^{t-l}P_{2}^{l})-P_{2}^{t}$
$=\sum_{l=1}^{t}(P_{1}^{t-l+1}P_{2}^{l-1}-P_{1}^{t-l}P_{2}^{l})$
$=\sum_{l=1}^{t}(P_{1}^{t-l}(P_{1}-P_{2})P_{2}^{l-1}).$
By applying the norm operator we find that:
$\|P_{1}^{t}-P_{2}^{t}\|_{2}$ $\leq\sum_{l=1}^{t}\|P_{1}^{t-l}\|_{2}\|P_{1}-P_{2}\|_{2}\|P_{2}^{l-1}\|_{2}$
$\leq tA^{N}\|P_{1}-P_{2}\|_{2},$
since $\|P\|_{2}\leq\sqrt{A^{N}}$ holds for all transition matrices $P$ over $\mathcal{A}^{N}$, and in particular for $P_{1}^{t-l}$ and $P_{2}^{l-1}$. By plugging the above into inequality (31) we find:
$\|\mu_{1}-\mu_{2}\|_{2}$ $\leq 2\sqrt{A^{N}}\|P_{1}^{t}-M\|_{TV}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+tA^{N}\|P_{1}-P_{2}\|_{2}.$
Finally, by choosing $t=t_{\operatorname{mix}}\left(1/\sqrt{16A^{N}}\right)$ we find:
$\|\mu_{1}-\mu_{2}\|_{2}\leq 2t_{\operatorname{mix}}\left(1/\sqrt{16A^{N}}\right)A^{N}\|P_{1}-P_{2}\|_{2}.$
The proof is then concluded by using the mixing-time bound from inequality (2).
Appendix E Numerical Illustration
We illustrate our convergence time results for log-linear learning for two-player identical interest games. In our setup, each player has the action set $\mathcal{A}=\{1,2,\ldots,10\}$, and the players have the same utility matrix denoted by $\{U(a_{1},a_{2})\}_{a_{1},a_{2}\in[10]}$. We simulate $10$ different utility matrices $U(\cdot,\cdot)\in[0,1]^{10\times 10}$ of the following form: we first fix a suboptimality gap $\Delta$, then we set $U(1,1)=1$, $U(10,10)=1-\Delta$, and the remaining entries of the utility matrix are sampled uniformly from the range $[0,1-\Delta)$.
Figure 1: Expected potential value when all players follow log-linear learning with $\beta$ set as in Equation (6). Left: performance for $\epsilon=0.05$ and different values of $\Delta$. Right: performance for $\Delta=0.1$ and different values $\epsilon$. Lines are averages over 10 randomly generated games, shaded areas represent one standard deviation, and the stars mark the first time the desired precision $1-\epsilon$ is reached.
In our first experiment, we fix $\epsilon$ as $0.05$ and vary the suboptimality gap $\Delta$ in $[0.15,0.10,0.075]$. We plot the expected potential value as a function of time when all players follow log-linear learning with temperature parameter $\beta$ set as in Equation (6). Figure 1 (left) demonstrates that the convergence time increases as the suboptimality gap $\Delta$ decreases. This shows that the suboptimality gap $\Delta$ quantifies the convergence time of log-linear learning well and therefore should be taken into account when bounding the convergence time, as done in our Theorem 3.1. In addition, we observe that this increase in convergence time is super-linear which aligns with the polynomial convergence time of $\tilde{\mathcal{O}}((\frac{1}{\epsilon})^{1/\Delta})$ that we proved in Theorem 3.1 when $\epsilon$ is smaller than $\Delta$.
In our second experiment, we fix $\Delta$ as $0.1$ and vary the precision level $\epsilon$ in $[0.1,0.05,0.025,0.02]$. We then plot the expected potential value as a function of time when all players follow log-linear learning with temperature parameter $\beta$ set as in Equation (6). Figure 1 (right) shows that the convergence time increases as the value of $\epsilon$ decreases. In other words, the convergence time of log-linear learning depends inversely on the desired suboptimality. This is also reflected in our convergence time bounds of order $\tilde{\mathcal{O}}((\frac{1}{\epsilon})^{1/\max\{\epsilon,\Delta\}})$ in Theorem 3.1.